OpenStudy (anonymous):

simplify: (2x(x+6)^2 - x^2(4)(x+6)^3)/(x+6)^8

OpenStudy (triciaal):

first work with the numerator and inside parentheses

OpenStudy (anonymous):

so i would pull out (x+6)^3?

OpenStudy (anonymous):

We have : You can divided same sentences and divided each on (x+6)^8 we have :(x+6)^2 and (x+6)^3 for divided .;)

OpenStudy (anonymous):

so its (x+6)^2 numerator (x+6)^3 denominator?

OpenStudy (anonymous):

would it just be (x+6)?

OpenStudy (anonymous):

No ! I think !:) Do you have a way ? Can you write you way ? @marissamtz

OpenStudy (anonymous):

i don't know what you mean!

OpenStudy (anonymous):

I get : 2x(1-2x)/(x+6)^3 @triciaal : Are you sure ???! Your denominator should change !

OpenStudy (anonymous):

Wait ...!

OpenStudy (anonymous):

I tried entering your answer and its wrong @E.ali but idk why!

OpenStudy (anonymous):

we have : (2x(x+6)^2 - x^2(4)(x+6)^3)/(x+6)^8=(2x(x+6)^2 - 4x^2(x+6)^2(x+6))/(x+6)^8= =((2x- 4x^2(x+6))(x+6)2)/(x+6)^8=2x-4x^2(x+6)/(x+6)^6=2x(1-2x(x+6)/(x+6)^6 I think answer is 2x(1-2x(x+6)/(x+6)^6

OpenStudy (campbell_st):

for me... I'd split the fraction \[\frac{2x(x +6)^2}{(x + 6)^8} - \frac{4x^2(x+6)^3}{(x+6)^8}\] simplify by removing common factors. which results in \[\frac{2x}{(x+6)^6} - \frac{4x^2}{(x + 6)^5}\] you could put it over a common denominator by \[\frac{2x}{(x+6)^6} - \frac{4x^2}{(x + 6)^5} \times \frac{(x + 6)}{(x + 6} = \frac{2x - 4x^2(x + 6)}{(x + 6)^6}\] and it could be simplified further if necessary

OpenStudy (anonymous):

Yea ! it s right ! you can factor 2x in : 2x-4x^2(x+6) in last ;)

OpenStudy (triciaal):

factor out (x+6)^2 (x+6)^2[(2x-4x^2(x+6)] (x+6)^2(2x-4x^3 +24x) denominator (x+6)^8 simplify by dividing by (x+6)^2 (2x-4x^3-24x^2)/(x + 6)^6

OpenStudy (anonymous):

!!! Right !